A student walks 625m west and 475m north, calculate the distance of the student walk. Calculate the displacement of student walking west.

Calculate the displacement of student walking north



1
Expert's answer
2022-03-28T16:13:49-0400

"d=x+y=1100~m,"

"r=\sqrt{x^2+y^2}=785~m."